# On computability of equilibrium states

### v2

@@ 4.3. Two subsets of $\mathcal{P}(X)$
namely, $\mu\in W^{c}.$ Hence, the claim holds. Now we return to the proof of Lemma 4.13. By Definition 3.21, Theorem 4.7, and Proposition 4.8, $\bigl{\{}(\varphi_{j}\cdot g_{k,n})\circ T^{-1}-\varphi_{j}\cdot g_{k,n}\cdot J_{\varphi}:j,\,k,\,n\in\mathbb{N}\text{ with }k\leqslant m\bigr{\}}$ is a uniformly computable sequence of functions. Then by Corollary 3.20, the family $\big{\{}\mathcal{I}_{j}^{k,n}:j,\,k,\,n\in\mathbb{N}\text{ with }k\leqslant m\big{\}}$ of integral operators $\mathcal{I}_{j}^{k,n}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ defined by
$\mathcal{I}_{j}^{k,n}(\mu)\coloneqq\int\!\bigl{(}(\varphi_{j}\cdot g_{k,n})\circ T^{-1}-\varphi_{j}\cdot g_{k,n}\cdot J_{\varphi}\bigr{)}\,\mathrm{d}\mu,\qquad\mu\in\mathcal{P}(X),$
is a uniformly computable sequence of functions. By Definition 3.5, $\mathbb{R}\setminus\{0\}$ is a lower-computable open set. Note that $\Phi_{j}^{k,n}=\bigl{(}\mathcal{I}_{j}^{k,n}\bigr{)}^{-1}(\mathbb{R}\setminus\{0\})$. Then by Corollary 3.11, $\Phi_{j}^{k,n}:j,\,k,\,n\in\mathbb{N}\text{ with }k\leqslant m$ is a uniformly computable sequence of lower-computable open sets. Therefore, by Proposition 3.7 and the claim, $W$ is a lower-computable open set. ∎
@@
We are now ready to complete the proof of Theorem 1.1.
@@
Assume that $X$ is recursively compact, then by Proposition 3.18, $\mathcal{P}(X)$ is recursively compact. By Corollary 2.9, there exists a unique $T$-invariant Gibbs state $\mu_{\varphi}$ for $T$ and $\varphi$, hence by Proposition 2.5, an unique equilibrium state $\mu_{\varphi}$ of $T$ for $\varphi$. Together with the definitions of $U$ and $W$, we have $\{\mu_{\varphi}\}=\mathcal{P}(X)\setminus(U\cup W)$. By Lemmas 4.12 and 4.13, $U$ and $W$ are both lower-computable open sets. It follows from Proposition 3.15 (ii) that $\{\mu_{\varphi}\}$ is recursively compact. Therefore, by Proposition 3.15 (i), the equilibrium state $\mu_{\varphi}$ of $T$ for $\varphi$ is computable. ∎
Applying Theorem 1.1, we establish Theorem 1.2.
@@
In the computable metric space $(D,\,\rho,\,\mathcal{S}^{\prime})$, by Theorem 1.1, there exists an algorithm which, on input $n\in\mathbb{N}$, outputs $\{k_{i}\}_{i=1}^{m}\subseteq\mathbb{Q}^{+}$ and $\{x_{i}\}_{i=1}^{m}\subseteq\mathcal{S}^{\prime}$ satisfying that $\sum_{i=1}^{m}k_{i}=1$, and $W_{\rho}(\mu_{\varphi},\mu_{n})\leqslant 2^{-n},$ where $\mu_{n}=\sum_{i=1}^{m}k_{i}\delta_{x_{i}}\in\mathcal{R}_{\mathcal{S}^{\prime}}$. Note that $\mathcal{S}^{\prime}$ is a uniformly computable sequence of points in the computable metric space $(X,\,\rho,\,\mathcal{S})$. So for each integer $1\leqslant i\leqslant m$, we can find a point $y_{i}\in\mathcal{S}$ such that $\rho\bigl{(}x_{i},y_{i}\bigr{)}<2^{-n}$. Define $\widehat{\mu}_{n}\coloneqq\sum_{i=1}^{m}k_{i}\delta_{y_{i}}\in\mathcal{R}_{\mathcal{S}}.$ Then
@@
Therefore, $\mu_{\varphi}$ is a computable point in the computable metric space $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{R}_{\mathcal{S}})$. ∎
@@ 5. Hyperbolic rational maps
In this section, we review some notations and results in complex dynamics and apply Theorem 1.2 to prove Theorem C. Recall that the spherical metric $d_{\widehat{\mathbb{C}}}$ on the Riemann sphere ${}^\widehat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, given by the length element $\mathrm{d}s(z)\coloneqq 2\lvert\mathrm{d}z\rvert/(1+\lvert z\rvert^{2})$ for each $z\in\widehat{\mathbb{C}}$, is a conformal metric. Moreover, (^$\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$ is a computable metric space, where $\mathbb{Q}^{2}$ is defined by $\{a+bi:a,\,b\in\mathbb{Q}\}$, and $\widehat{\mathbb{C}}$ is recursively compact. The spherical derivative $f^{\#}$ of a holomorphic function $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is given by
(5.1) $f^{\#}(z)\coloneqq\lim\limits_{w\to z}\frac{d_{\widehat{\mathbb{C}}}(f(w),f(z))}{d_{\widehat{\mathbb{C}}}(w,z)}=\frac{1+\lvert z\rvert^{2}}{1+\lvert f(z)\rvert^{2}}\cdot\lvert f^{\prime}(z)\rvert$
for each $z\in\widehat{\mathbb{C}}$. If $z$ or $f(z)$ is equal to $\infty$, then the last expression of (5.1) has to be understood as a suitable limit. We identify $\widehat{\mathbb{C}}$ with the unit sphere in $\mathbb{R}^{3}$ via stereographic projection. The chordal metric $\sigma$ on $\widehat{\mathbb{C}}$ is the metric that corresponds to the Euclidean metric in $\mathbb{R}^{3}$ under this identification. More precisely, we have
(5.2) $\sigma(z,w)=\frac{2\lvert z-w\rvert}{\sqrt{1+\lvert z\rvert^{2}}\sqrt{1+\lvert w\rvert^{2}}}\quad\text{for each }z,\,w\in\mathbb{C},$
and $\sigma(\infty,z)=\sigma(z,\infty)=2\big{/}\sqrt{1+\lvert z\rvert^{2}}$ for each $z\in\mathbb{C}$. Any holomorphic map $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ on the Riemann sphere can be expressed as a rational function, that is, as the quotient $f(z)=p(z)/{q(z)}$ of two polynomials. Here we may assume that $p(z)$ and $q(z)$ have no common roots. The degree $d$ of the rational map $f$ is defined as the maximum of the degrees of $p$ and $q$. The Fatou set $\mathcal{F}_{f}$ of $f$ is defined to be the set of points $z\in\widehat{\mathbb{C}}$ such that there exists an open neighborhood $U(z)$ on which the family $\{f^{n}|_{U(z)}\}_{n\in\mathbb{N}}$ is equicontinuous with respect to the spherical metric $d_{\widehat{\mathbb{C}}}$. The Julia set $\mathcal{J}_{f}$ of $f$ is the complement of the Fatou set $\mathcal{F}_{f}$.
@@
If $-^f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is a rational map whose degree $d\geqslant 2$, then for each $z_{0}\in\mathcal{J}_{f}$, the set $\bigcup_{n\in\mathbb{N}}f^{-n}(z_{0})$ of all iterated preimages of $z_{0}$ is dense in $\mathcal{J}_{f}$.
The map we consider in this section is a hyperbolic rational map, which is defined as follows (see for example, [Mil06, Theorem 19.1]).
@@
Lemma 5.3 ([CG93, Lemma 2.1]).
Assume that $-^f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is a hyperbolic rational map whose degree is at least two with $\infty\notin\mathcal{J}_{f}$. Then there exist two constants $a>0$ and $A>1$ for all $n\in\mathbb{N}$ and $z\in\mathcal{J}_{f}$, $\lvert(f^{n})^{\prime}(z)\rvert\geqslant aA^{n}$.
Lemma 5.4 ([CG93, Theorem 2.3]).
If $f$ is a hyperbolic rational map whose degree is at least two, then $\mathcal{J}_{f}$ has zero area.
The following lemma is due independently to Braverman [Br04] (in the case of polynomials) and Rettinger [Re04] (in the case of rational maps).
@@
If $f$ is a computable hyperbolic rational map whose degree is at least two, then the distance function $d_{\widehat{\mathbb{C}}}(\mathcal{J}_{f},x)$ is computable.
@@
Let $-^f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a computable hyperbolic rational map of degree $d\geqslant 2$. Then for each repelling periodic point $w$ of $f$, the following statements hold:
@@
    There is an enumeration $\{w_{i}\}_{i}$ of the set $\bigcup_{n\in\mathbb{N}}f^{-n}(w)$ such that $\{w_{i}\}_{i}$ is a uniformly computable sequence of points in the computable metric space (^$\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$;
@@
(i) First, we demonstrate that every periodic point of $f$ is computable in the computable metric space (^$\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$. Let $u$ be a such point, then $f^{m}(u)=u$ for some $m\in\mathbb{N}$. To establish this, we apply a standard root-finding algorithm for the function $f^{m}(x)-x$ (see for example, [BBY10, Appendix A]) to compute $u$. Hence, $w$ is computable. By [Mil06, Lemma 4.6], we obtain that $w\in\mathcal{J}_{f}$. Hence, by the same root-finding algorithm, we can compute all the roots of the function $g_{n}(x)=f^{n}(x)-w$ for each $n\in\mathbb{N}$ and enumerate them as $\{w_{i}\}_{i}$. Hence, $\{w_{i}\}_{i}$ is a uniformly computable sequence of points in the computable metric space (^$\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$. (ii) By Proposition 5.1, $\bigcup_{n\in\mathbb{N}}f^{-n}(w)$ is dense in $\mathcal{J}_{f}$. Together with statement (i), one can see that $\bigl{(}\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$ is a computable metric space. Hence, to demonstrate statement (ii), by Proposition 4 in [GHR10], it suffices to prove that $\bigl{(}\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$ is recursively precompact. Fix an integer $m$. Since (^$\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$ is recursively precompact, we can compute a $2^{-m-1}$-net $N^{\prime\prime}_{m}$ of $\widehat{\mathbb{C}}$. By Lemma 5.5, we can compute a $2^{-n-1}$-net $N^{\prime}_{m}\subseteq N^{\prime\prime}_{m}$ of $\mathcal{J}_{f}$ satisfying that $\mathcal{J}_{f}\cap B(y,2^{-m-1})\neq\emptyset$ for each $y\in N^{\prime}_{m}$. By Proposition 5.1 and statement (i), we can compute a point $t(y)\in B(y,2^{-m-1})\cap\bigl{(}\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$ for each $y\in N^{\prime}_{m}$. Hence, the set $\{t(y):y\in N^{\prime}_{m}\}\subseteq\mathcal{J}_{f}$ is a $2^{-m}$-net of the Julia set $\mathcal{J}_{f}$. This completes the proof. ∎
@@
Let $-^f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a hyperbolic rational map of degree $d\geqslant 2$ with $\infty\notin\mathcal{J}_{f}$. Let $C_{f}\coloneqq f(\{z\in\widehat{\mathbb{C}}:f^{\prime}(z)=0\})$ be the $f$-image of the critical set of $f$. Assume that there exist constants $V_{1}$, $V_{2}$, $V_{3}$, and $V_{4}$ satisfying the following properties for each pair $x,\,y\in\mathcal{J}_{f}$:
@@
Then the following statements hold:
@@
    $f$ is injective on $B(x,l)$ for each $x\in\mathcal{J}_{f}$, where $l\coloneqq V_{1}/(4V_{2})$.
@@
    $f$ is distance-expanding on $\mathcal{J}_{f}$ with respect to the spherical metric with constants $\eta\coloneqq V_{4}l$, $\lambda\coloneqq V_{2}V_{4}/8$, and $\xi\coloneqq V_{2}l/8$, where $l$ is from statement (i).
@@
(i) Fix an arbitrary point $x\in\mathcal{J}_{f}$. By property (1), one can see that there exists no critical value in $B(f(x),V_{1})$. Hence, we can define a conformal inverse $f^{-1}_{x}\colon B(f(x),V_{1})\rightarrow\widehat{\mathbb{C}}$ in the sense of $f(f^{-1}_{x}(y))=y$ for each $y\in B(f(x),V_{1})$ and $f^{-1}_{x}(f(x))=x$. Then by Köbe’s distorsion theorem, we have $B(x,\lvert(f_{x}^{-1})^{\prime}(f(x))\rvert V_{1}/4)\subseteq f^{-1}_{x}(B(f(x),V_{1}))$. By the Chain Rule and property (2), $\lvert(f_{x}^{-1})^{\prime}(f(x))\rvert=1/\lvert f^{\prime}(x)\rvert\geqslant V_{3}^{-1}$. Thus, one can see that $l=V_{1}/(4V_{3})\leqslant\lvert(f_{x}^{-1})^{\prime}(f(x))\rvert V_{1}/4$. Therefore, $f$ is injective on $B(x,l)$. (ii) Fix an arbitrary pair of $x,\,y\in\mathcal{J}_{f}$ with $d_{\widehat{\mathbb{C}}}(x,y)\leqslant\eta$. First, we demonstrate that $d_{\widehat{\mathbb{C}}}(f(x),f(y))\geqslant\lambda d_{\widehat{\mathbb{C}}}(x,y)$. By property (3), we obtain that $\lvert x-y\rvert\leqslant\eta/V_{4}=l$. Hence, by statement (i), $f$ is injective on $B(x,\lvert x-y\rvert)$. By Köbe’s distorsion theorem, we have
$B(f(x),\lvert x-y\rvert\cdot\lvert f^{\prime}(x)\rvert/4)\subseteq f(B(x,\lvert x-y\rvert)).$
Noting that $f(y)\in\partial f(B(x,\lvert x-y\rvert))$, one can see that $\lvert f(x)-f(y)\rvert\geqslant\lvert x-y\rvert\cdot\lvert f^{\prime}(x)\rvert/4\geqslant V_{2}\lvert x-y\rvert/4$. Since $\mathcal{J}_{f}$ is $f$-invariant, then $f(x),\,f(y)\in\mathcal{J}_{f}$. By property (3), we have

### v3

@@ 4.3. Two subsets of $\mathcal{P}(X)$
namely, $\mu\in W^{c}.$ Hence, the claim holds.
By Definition 3.21, Theorem 4.7, and Proposition 4.8,
$\bigl{\{}(\varphi_{j}\cdot g_{k,n})\circ T^{-1}-\varphi_{j}\cdot g_{k,n}\cdot J_{\varphi}:j,\,k,\,n\in\mathbb{N}\text{ with }k\leqslant m\bigr{\}}$
is a uniformly computable sequence of functions. Then by Corollary 3.20, the family $\mathcal{I}_{j}^{k,n}\colon\mathcal{P}(X)\rightarrow\mathbb{R},\,j,\,k,\,n\in\mathbb{N}\text{ with }k\leqslant m,$ defined by
$\mathcal{I}_{j}^{k,n}(\mu)\coloneqq\int\!\bigl{(}(\varphi_{j}\cdot g_{k,n})\circ T^{-1}-\varphi_{j}\cdot g_{k,n}\cdot J_{\varphi}\bigr{)}\,\mathrm{d}\mu\quad\text{ for all }\mu\in\mathcal{P}(X)\text{ and }j,\,k,\,n\in\mathbb{N}\text{ with }k\leqslant m,$
is a uniformly computable sequence of functions. By Definition 3.5, $\mathbb{R}\smallsetminus\{0\}$ is a lower-computable open set. Note that $\Phi_{j}^{k,n}=\bigl{(}\mathcal{I}_{j}^{k,n}\bigr{)}^{-1}(\mathbb{R}\smallsetminus\{0\})$. Then by Corollary 3.11, $\Phi_{j}^{k,n},\,j,\,k,\,n\in\mathbb{N}\text{ with }k\leqslant m$ form a uniformly computable sequence of lower-computable open sets. Therefore, by Proposition 3.7 and the claim, $W$ is a lower-computable open set. ∎
@@
Now we prove Theorem 1.1.
@@
Since $X$ is recursively compact, by Proposition 3.18, $\mathcal{P}(X)$ is also recursively compact. By Corollary 2.9, there exists a unique $T$-invariant Gibbs state $\mu_{\varphi}$ for $T$ and $\varphi$. By Proposition 2.6, $\mu_{\varphi}$ is also a unique equilibrium state $\mu_{\varphi}$ for $T$ and $\varphi$. Together with the definitions of $U$ and $W$, we have $\{\mu_{\varphi}\}=\mathcal{P}(X)\smallsetminus(U\cup W)$. By Lemmas 4.12 and 4.13, $U$ and $W$ are both lower-computable open sets. It follows from Proposition 3.15 (iii) that $\{\mu_{\varphi}\}$ is recursively compact. Therefore, by Proposition 3.15 (i), the measure $\mu_{\varphi}$ is computable. ∎
Next, we apply Theorem 1.1 to establish Theorem 1.2 as follows.
@@
In the computable metric space $(D,\,\rho^{\prime},\,\mathcal{S}^{\prime})$, by Theorem 1.1, there exists an algorithm which, on input $n\in\mathbb{N}$, outputs finite sequences $\{k_{i}\}_{i=1}^{m}$ in $\mathbb{Q}^{+}$ and $\{x_{i}\}_{i=1}^{m}$ in $\mathcal{S}^{\prime}$ satisfying $\sum_{i=1}^{m}k_{i}=1$, and $W_{\rho^{\prime}}(\mu_{\varphi},\mu_{n})\leqslant 2^{-n},$ where $\mu_{n}=\sum_{i=1}^{m}k_{i}\delta_{x_{i}}\in\mathcal{R}_{\mathcal{S}^{\prime}}$. Since $\rho^{\prime}$ is the restriction of the metric $\rho$, we have $W_{\rho}(\mu_{\varphi},\mu_{n})\leqslant 2^{-n}$. Note that $\mathcal{S}^{\prime}$ is a uniformly computable sequence of points in the computable metric space $(X,\,\rho,\,\mathcal{S})$. Then for each integer $i\in[1,m]$, we can find a point $y_{i}\in\mathcal{S}$ such that $\rho\bigl{(}x_{i},y_{i}\bigr{)}<2^{-n}$. Define $\widehat{\mu}_{n}\coloneqq\sum_{i=1}^{m}k_{i}\delta_{y_{i}}\in\mathcal{R}_{\mathcal{S}}.$ Thus we have $W_{\rho}(\mu_{n},\widehat{\mu}_{n})<2^{-n}$. Then
@@
Hence, $\mu_{\varphi}$ is a computable point in the computable metric space $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{R}_{\mathcal{S}})$. ∎
@@ 5. Hyperbolic rational maps
In this section, we review some notations and results in complex dynamics and apply Theorem 1.2 to prove Theorem C.
Recall that the spherical metric $d_{\widehat{\mathbb{C}}}$ on the Riemann sphere $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$, given by the length element $\mathrm{d}s(z)\coloneqq 2\lvert\mathrm{d}z\rvert/(1+\lvert z\rvert^{2})$ for each $z\in\widehat{\mathbb{C}}$, is a conformal metric. Moreover, $\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$ is a computable metric space and $\widehat{\mathbb{C}}$ is recursively compact, where $\mathbb{Q}^{2}\coloneqq\{a+bi:a,\,b\in\mathbb{Q}\}$. The spherical derivative $f^{\#}$ of a holomorphic function $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is given by
(5.1) $f^{\#}(z)\coloneqq\lim\limits_{w\to z}d_{\widehat{\mathbb{C}}}(f(w),f(z))\big{/}d_{\widehat{\mathbb{C}}}(w,z)=\lvert f^{\prime}(z)\rvert\cdot\bigl{(}1+\lvert z\rvert^{2}\bigr{)}\big{/}\bigl{(}1+\lvert f(z)\rvert^{2}\bigr{)}$
for each $z\in\widehat{\mathbb{C}}$. If $z$ or $f(z)$ is equal to $\infty$, then the last expression of (5.1) has to be understood as a suitable limit.
We identify $\widehat{\mathbb{C}}$ with the unit sphere in $\mathbb{R}^{3}$ via stereographic projection. The chordal metric $\sigma$ on $\widehat{\mathbb{C}}$ is the metric that corresponds to the Euclidean metric in $\mathbb{R}^{3}$ under this identification. More precisely,
(5.2) $\sigma(z,w)=2\lvert z-w\rvert\big{/}\sqrt{(1+\lvert z\rvert^{2})(1+\lvert w\rvert^{2})}\quad\text{ for each pair of }z,\,w\in\mathbb{C},$
and $\sigma(\infty,z)=\sigma(z,\infty)=2\big{/}\sqrt{1+\lvert z\rvert^{2}}$ for each $z\in\mathbb{C}$.
Any holomorphic map $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ on the Riemann sphere can be expressed as a rational function, that is, as the quotient f(z)=p(z)/q(z)f(z)=p(z)/{q(z)} of two polynomials $p$ and $q$. Here we may assume that $p$ and $q$ have no common roots. The degree $d$ of the rational map $f$ is defined as the maximum of the degrees of $p$ and $q$. The Fatou set $\mathcal{F}_{f}$ of $f$ is defined to be the set of points $z\in\widehat{\mathbb{C}}$ satisfying that there exists an open neighborhood $U(z)$ of $z$ on which the family $\{f^{n}|_{U(z)}\}_{n\in\mathbb{N}}$ is equicontinuous with respect to the spherical metric $d_{\widehat{\mathbb{C}}}$. The Julia set $\mathcal{J}_{f}$ of $f$ is the complement of the Fatou set $\mathcal{F}_{f}$.
@@
If $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is a rational map of degree $d\geqslant 2$, then for each $z_{0}\in\mathcal{J}_{f}$, the set $\bigcup_{n\in\mathbb{N}}f^{-n}(z_{0})$ of all iterated preimages of $z_{0}$ is dense in $\mathcal{J}_{f}$.
The maps we consider in this section are hyperbolic rational maps, which are defined as follows (see e.g. [Mil06, Theorem 19.1]).
@@
Lemma 5.3 ([CG93, Lemma V.2.1]).
Assume that $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is a hyperbolic rational map of degree $d\geqslant 2$ with $\infty\notin\mathcal{J}_{f}$. Then there exists $a>0$ and $A>1$ such that $\lvert(f^{n})^{\prime}(z)\rvert\geqslant aA^{n}$ for all $n\in\mathbb{N}$ and $z\in\mathcal{J}_{f}$.
Lemma 5.4 ([CG93, Theorem V.2.3]).
If $f$ is a hyperbolic rational map of degree $d\geqslant 2$, then $\mathcal{J}_{f}$ has zero area.
The following lemma is due independently to Braverman [Br04] (in the case of the hyperbolic polynomials) and Rettinger [Re05] (in the case of the hyperbolic rational maps).
@@
If $f$ is a computable hyperbolic rational map of degree $d\geqslant 2$, then the distance function $x\mapsto d_{\widehat{\mathbb{C}}}(\mathcal{J}_{f},x)$ is computable.
@@
Let $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a computable hyperbolic rational map of degree $d\geqslant 2$. Then for each repelling periodic point $w$ of $f$, the following statements are true:
@@
    There is an enumeration $\{w_{i}\}_{i\in\mathbb{N}}$ of the set $\bigcup_{n\in\mathbb{N}}f^{-n}(w)$ such that $\{w_{i}\}_{i\in\mathbb{N}}$ is a uniformly computable sequence of points in the computable metric space $\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$.
@@
(i) First, we demonstrate that $w$ is computable in the computable metric space $\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$. Since $w$ is a periodic point of $f$, there exists $m\in\mathbb{N}$ with $f^{m}(w)=w$. Then we apply a standard root-finding algorithm (see e.g. [BBY12, Appendix A]) for the function $f^{m}(x)-x$ to compute $w$. Thus $w$ is computable and lies in $\mathcal{J}_{f}$ by [Mil06, Lemma 4.6]. Hence, by the same root-finding algorithm, we can compute all the roots of the function $g_{n}(x)=f^{n}(x)-w$ for each $n\in\mathbb{N}$. Hence, there is an enumeration $\{w_{i}\}_{i\in\mathbb{N}}$ of the set $\bigcup_{n\in\mathbb{N}}f^{-n}(w)$ such that $\{w_{i}\}_{i\in\mathbb{N}}$ is a uniformly computable sequence of points in the computable metric space $\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$.
(ii) By Proposition 5.1, $\bigcup_{n\in\mathbb{N}}f^{-n}(w)$ is dense in $\mathcal{J}_{f}$. Together with statement (i), it is not hard to see that $\bigl{(}\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$ is a computable metric space. Hence, to demonstrate statement (ii), by Proposition 4 in [GHR11], it suffices to prove that $\bigl{(}\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$ is recursively precompact.
Now we fix an integer $s$ and construct a $2^{-s}$-net of the Julia set $\mathcal{J}_{f}$ as follows. Since $\bigl{(}\widehat{\mathbb{C}},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2}\bigr{)}$ is recursively precompact, we can compute a $2^{-s-1}$-net $N^{\prime\prime}_{s}$ of $\widehat{\mathbb{C}}$. By Lemma 5.5, we can compute a $2^{-s-1}$-net $N^{\prime}_{s}\subseteq N^{\prime\prime}_{s}$ of $\mathcal{J}_{f}$ satisfying $\mathcal{J}_{f}\cap B\bigl{(}y,2^{-s-1}\bigr{)}\neq\emptyset$ for each $y\in N^{\prime}_{s}$. By Proposition 5.1 and statement (i), we can compute a point $t(y)\in B\bigl{(}y,2^{-s-1}\bigr{)}\cap\bigl{(}\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$ for each $y\in N^{\prime}_{s}$. Hence, $\{t(y):y\in N^{\prime}_{s}\}\subseteq\mathcal{J}_{f}$ is a $2^{-s}$-net of the Julia set $\mathcal{J}_{f}$. This completes the proof. ∎
@@
Let $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a hyperbolic rational map of degree $d\geqslant 2$ with $\infty\notin\mathcal{J}_{f}$. Let $C_{f}\coloneqq f(\{z\in\widehat{\mathbb{C}}:f^{\prime}(z)=0\})$ be the set of critical values of $f$. Assume that there exist positive constants $V_{1}$, $V_{2}$, $V_{3}$, and $V_{4}$ satisfying the following properties for each pair of $x,\,y\in\mathcal{J}_{f}$:
@@
Put $l\coloneqq V_{1}/(4V_{3})$. Then the following statements are true:
@@
    The map $f$ is injective on $B(x,l)$ for each $x\in\mathcal{J}_{f}$.
@@
    The restriction $f|_{\mathcal{J}_{f}}\colon\mathcal{J}_{f}\rightarrow\mathcal{J}_{f}$ is a distance-expanding map with respect to $d_{\widehat{\mathbb{C}}}$ with constants $\eta\coloneqq V_{4}l/2$, $\lambda\coloneqq V_{2}V_{4}/8$, and $\xi\coloneqq V_{2}l/16$.
@@
(i) Consider an arbitrary point $x\in\mathcal{J}_{f}$. Then $f(x)\in\mathcal{J}_{f}$. By property (1), we obtain that $B(f(x),V_{1})\cap C_{f}=\emptyset$. Hence, we can define a conformal inverse $f^{-1}_{x}\colon B(f(x),V_{1})\rightarrow\widehat{\mathbb{C}}$ such that $f^{-1}_{x}(f(x))=x$ and $f\bigl{(}f^{-1}_{x}(y)\bigr{)}=y$ for each $y\in B(f(x),V_{1})$. Then by Koebe’s one-quarter theorem (see e.g. [CG93, Theorem I.1.3]), we have $B(x,\lvert(f_{x}^{-1})^{\prime}(f(x))\rvert V_{1}/4)\subseteq f^{-1}_{x}(B(f(x),V_{1}))$. By the chain rule and property (2), $\lvert(f_{x}^{-1})^{\prime}(f(x))\rvert=1/\lvert f^{\prime}(x)\rvert\geqslant V_{3}^{-1}$. Thus, one can see that $l=V_{1}/(4V_{3})\leqslant\lvert(f_{x}^{-1})^{\prime}(f(x))\rvert V_{1}/4$. Therefore, $f$ is injective on $B(x,l)$.
[…]
$).B(f(x),\lvert x-y\rvert\cdot\lvert f^{\prime}(x)\rvert/4)\subseteq f(B(x,\lvert x-y\rvert$)).
By $)f(y)\in\partial f(B(x,\lvert x-y\rvert$)), we obtain that $\lvert f(x)-f(y)\rvert\geqslant\lvert x-y\rvert\cdot\lvert f^{\prime}(x)\rvert/4\geqslant V_{2}\lvert x-y\rvert/4$. Note that by [CG93, Theorem III.1.3], $f(\mathcal{J}_{f})=\mathcal{J}_{f}$. Then $f(x)$ and $f(y)$ both still belong to $\mathcal{J}_{f}$. By property (3), we have